Show that two events A and B are independent if and only if P(A ∩ B) = P(A)P(B) when
at least one of P(A) or P(B) is not zero.
Thank you.

Answer by Edwin McCravy(20086)

Many textbooks define independent events by that formula.  You can't prove
what is stated in a definition.

Two events A, B are said to be independent if and only if 

P(A ∩ B) = P(A)P(B).

All I can guess is that your textbook has defined it this way:

Two events A, B are said to be independent if and only if 

P(A|B) = P(A)

And the definition of conditional probability is 

P(A|B) = P(A ∩ B)/P(B).

If so, you can prove it by substituting P(A) for P(A|B)

P(A) = P(A ∩ B)/P(B)

and then multiplying both sides by P(B)

P(A)P(B) = P(A ∩ B)

Edwin